# Information About Other Agents in Mechanism Design

### v2

@@ 3 Partial implementation and the direct mechanism
Given any SCF $F$ we can consider the direct mechanism defined by $\Gamma_{F}=\langle M,h\rangle,$ where $M_{i}=\Theta_{i}$ for each $i$ and $h=F.$ Each agent $i$ has a truth-telling strategy $s:\Theta_{i}\to\Theta_{i}$ given by $s(\theta_{i})=\theta_{i}.$ The revelation principle says that $F$ is partially implementable with respect to $\{\phi_{i}\}$ iff truth-telling for all agents is a Nash equilibrium of the direct mechanism. This is equivalent to incentive compatibility, which for the noninformative information structure $\{\phi_{i}\}$ means that for each agent $i\in I$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$,
@@
If, however, $\sigma_{i}\in\{A_{i1},\dots,A_{ik}\}$ contained information about $\Theta_{-i}$, partial implementation requires $F$ to be incentive-compatible with respect to the beliefs of agent $i$ in every realization of $\sigma_{i}.$ In other words, for every $1\leq\ell\leq k$ it must be that
$\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{\ell})u_{i}(F(\theta_{i},\theta_{-i}),\theta_{i})\geq\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{\ell})u_{i}(F(\theta_{i}^{\prime},\theta_{-i}),\theta_{i}).$
The law of total expectations implies that  (3) is just a weighted average of the equations  (3) for $1\leq\ell\leq k.$ This implies the following proposition.
@@ 4.1 Incentive compatibility
To see that $F$ is incentive-compatible, truth-telling is must be an equilibrium of the direct mechanism. For agent 1, observe
@@ 4.4 Full implementability of $F$
It follows that $\alpha_{1}$ is undermined by agent 1. Similar arguments show that $\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ are also undermined. Since $\sigma_{3}$ only affects the information of agent 3, each $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2 in both realizations of $\sigma_{3}$. In the previous subsection we saw that agent 3 can undermine $\alpha$ in either realization of $\sigma_{3}$. Therefore, lemma 1 implies that $F$ is fully implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$.
@@ Appendix B Appendix B: Omitted proofs
Fix a agent $j.$ Since every agents’ beliefs and every signal has full support, the only possible signal that affects agent $j$’s expected payoff is $\sigma_{j}.$ Suppose $\sigma_{j}\in\{A_{j1},\dots,A_{jk}\}$ and let $q_{\ell}$ be the unconditional probability that $\sigma_{j}=A_{\ell}$ (so $q_{\ell}$ is averaged over all possible states). If $F$ is incentive-compatible with respect to agent $j$ after agent $j$ receives this signal, then
$q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{\ell})\geq q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{\ell})$
@@
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{\ell})=\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j}),$
@@
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{\ell})=\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j}).$

### v3

@@ 3 Partial implementation and the direct mechanism
Given any SCF $F$ we can consider the direct mechanism defined by $\Gamma_{F}=\langle M,h\rangle,$ where $M_{i}=\Theta_{i}$ for each $i$ and $h=F.$ Each agent $i$ has a truth-telling strategy $s:\Theta_{i}\to\Theta_{i}$ given by $s(\theta_{i})=\theta_{i}.$ The revelation principle says that $F$ is partially implementable with respect to $\{\phi_{i}\}$ iff for each agent $i\in I$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$,
@@
If, however, $\sigma_{i}\in\{A_{i1},\dots,A_{ik}\}$ contained information about $\Theta_{-i}$, partial implementation requires $F$ to be incentive-compatible with respect to the beliefs of agent $i$ in every realization of $\sigma_{i}.$ In other words, for every $A_{i\ell}$ and $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ it must be that
$\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{i\ell})u_{i}(F(\theta_{i},\theta_{-i}),\theta_{i})\geq\sum_{\theta_{-i}\in\Theta_{-i}}\mu(\theta_{-i}\mid\theta_{i},\sigma_{i}=A_{i\ell})u_{i}(F(\theta_{i}^{\prime},\theta_{-i}),\theta_{i}).$
The law of total expectation implies that  (3) is just a weighted average of the equations  (3) for $1\leq\ell\leq k.$ This implies the following proposition.
@@ 4.1 Incentive compatibility
To see that $F$ is incentive-compatible, truth-telling must be an equilibrium of the direct mechanism. For agent 1, observe
@@ 4.4 Full implementability of $F$
It follows that $\alpha_{1}$ is undermined by agent 1. Similar arguments show that $\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ are also undermined. Since $\sigma_{3}$ only affects the information of agent 3, each $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2 in both realizations of $\sigma_{3}$. In the previous subsection we saw that agent 3 can undermine $\alpha$ in either realization. Therefore, lemma 1 implies that $F$ is fully implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$.
@@ Appendix B Appendix B: Omitted proofs
Fix a agent $j.$ Since every agents’ beliefs and every signal has full support, the only possible signal that affects agent $j$’s expected payoff is $\sigma_{j}.$ Suppose $\sigma_{j}\in\{A_{j1},\dots,A_{jk}\}$ and let $q_{\ell}$ be the unconditional probability that $\sigma_{j}=A_{j\ell}$ (so $q_{\ell}$ is averaged over all possible states). If $F$ is incentive-compatible with respect to agent $j$ after agent $j$ receives this signal, then
$q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})\geq q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})$
@@
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})=\mathbb{E}(u_{j}(F(\theta_{-j},\theta_{j}),\theta_{j})\mid\theta_{j}),$
@@
$\sum_{\ell=1}^{k}q_{\ell}\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j},\sigma_{j}=A_{j\ell})=\mathbb{E}(u_{j}(F(\theta_{-i},\theta_{j}^{\prime}),\theta_{j})\mid\theta_{j}).$